You own a $25,000 portfolio consisting of two stocks, A and B. Stock A is valued at $10,000 and has an expected return of 10%. Stock B has an expected return of 5%. What are the weights of stock A and B in the portfolio?
Answer #1
Value of stock B = Total portfolio value - Value of stock A = 25000 - 10000 15000
Weight of stock A = Value of stock A / Total portfolio value = 10000 / 25000 40%
Weight of stock B = Value of stock B / Total portfolio value = 15000 / 25000 60%